OpenStudy (johnnydicamillo):

a balloon ascending at a rte of 12ft/s is at a height of 80ft above the ground when a package is dropped. How long does it take the package to reach the ground?

OpenStudy (johnnydicamillo):

\[a(t) = -32\] \[v(t) = -32t + c\]

OpenStudy (johnnydicamillo):

the c is 12 so \[v(t) = -32t + 12\] \[s(t) = -16t^2 + 12t + d\] and d = 80 so \[s(t) = -16t^2 + 12t + 80\]

OpenStudy (johnnydicamillo):

This doesn't look like its going to reduce, nicely, I think I did something wrong. Also I forgot to mention that acceleration due to gravity is 32ft/s

OpenStudy (lilah):

@Hero

OpenStudy (lilah):

@TheSmartOne

OpenStudy (lilah):

@robtobey2

OpenStudy (lilah):

@Jadeishere

OpenStudy (jadeishere):

I'm not the best person to ask! Terribly sorry! @mathmate

OpenStudy (johnnydicamillo):

I think I just need to reduce that last part and solve for 0 but I am not sure how.

OpenStudy (johnnydicamillo):

quadratic formula

OpenStudy (johnnydicamillo):

\[ax^2 + bx + c = 0\] \[x = \frac{ -b +-\sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (faiqraees):

Just use the equation S=ut+ 1/2 a t^2 U = 12ft/s S = displacement = 80 ft a = -32ft/s^2 Find t

OpenStudy (faiqraees):

80 = 12t - 16t² 16t²-12t+80 = 0 Solve this

OpenStudy (johnnydicamillo):

right, that is the issue...

OpenStudy (johnnydicamillo):

okay, but I don't know this part: \[16^{2} + 12t = -80\]

OpenStudy (johnnydicamillo):

I think I could just use the quadratic formula

OpenStudy (johnnydicamillo):

\[t = \frac{ -12 +- \sqrt{12^2 - 4 (-16)(80)} }{ 2(-16) }\]

OpenStudy (faiqraees):

Are you sure the numbers are correct/?

OpenStudy (johnnydicamillo):

are you asking because you know the numbers are incorrect?

OpenStudy (faiqraees):

Well because the answer is coming in imaginary form

OpenStudy (johnnydicamillo):

yeah, I am pretty sure the numbers are correct, but the 16t²-12t+80 = 0 can reduce to 4t^2 - 3t + 20 = 0

OpenStudy (johnnydicamillo):

solved! \[-3+ \sqrt{\frac{ 329 }{ 8 }}\] verified with solution.

OpenStudy (johnnydicamillo):

also with wolframalpha: http://www.wolframalpha.com/input/?i=0+%3D+-16x%5E2+%2B+12x+%2B+80,+find+x

OpenStudy (faiqraees):

Oh yep got it

OpenStudy (johnnydicamillo):

because \[t = -3 + \sqrt{\frac{ -3^2 - 4(4)(20) }{ 2(4) }}\] goes to \[t = -3 + \sqrt{\frac{ -9 - 320}{ 8 }}\]

OpenStudy (mathmate):

If we use upwards as positive, the distance travelled is -80 feet, and the initial speed is +12 ft/s. This comes out to be \(16t^2-12t-80=0\) and the quadratic formula will give two real solutions. Reject the negative value. The solution retained should be between 2 and 3 seconds.